The half-life for the radioactive decay of U−238 is 4.5 billion years and is independent of initial concentration.

1.How long will it take for 10% of the U−238 atoms in a sample of U−238 to decay?

2.If a sample of U−238 initially contained 1.9×1018 atoms when the universe was formed 13.8 billion years ago, how many U−238 atoms will it contain today?
Answer #1

1. half life time(t1/2) = 4.5*10^9 years

K1 = 0.693/T1/2

     = 0.693/(4.5*10^9)

     = 1.54*10^-10 y-1

k = (1/t)ln(a0/a)

a0 = initial concentration = 100%

a = concentration remains after time t years = 90%

1.54*10^-10 = (1/t)ln(100/90)

t = 6.84*10^8 years

2.

k = (1/t)ln(a0/a)

a0 = initial concentration = 1.9*10^18 atoms

a = concentration remains after time t years = x

t = 13.8 billion = 13.8*10^9 years

1.54*10^-10 = (1/(13.8*10^9))ln((1.9*10^18)/x)

x = 2.27*10^17 atoms

U−238 atoms contain today = 2.27*10^7 atoms